# EvidenceBench: A Benchmark for Extracting Evidence from Biomedical Papers

### v1

@@ 1 Introduction
We conduct a benchmarking study on a variety of large language models (LLMs) and embedding models, which provide several insights. First, although LLMs still fall short of expert-level performance on this task, indicating that they cannot replace humans, they have the potential to assist them. Second, embedding models consistently underperform compared to large language models due to lack of global context from research papers. Third, we report that current LLMs trained for long document understanding still get "Lost in the Middle" (Liu et al., 2024). See Appendix J for details of our analyses.
@@ 2.1 Definitions
$\mathcal{S}(f_{i},s_{j})=\begin{cases}1&\text{if }s_{j}\text{ is a the source of information for }f_{i}\\ 0&\text{otherwise}\end{cases}$
@@ 3.1 Data Sources
There are two data sources for EvidenceBench and EvidenceBench-100k. First, a collection of 107,887 CC-BY open-sourced biomedical research papers where each research paper represents a datapoint. Second, a collection of 44,772 review papers from PubMed Central. Each biomedical research paper has a corresponding evidence summary included in a review paper. Specifically, EvidenceBench has 426 datapoints and EvidenceBench-100k has 107,461 datapoints.
@@ 3.2 Dataset Pipeline Overview
Harvesting the Evidence Summary: The highlighted portion in the review paper in Figure 2 is an evidence summary. Evidence summary has an XML citation embedded in it, so it can be identified and harvested by a deterministic algorithm. See Appendix G for details.
@@ 3.4 Alignment Annotation of Study Aspects and Sentences
Each team labeled 8111 (sentence, study aspect) pairs in total. Table 2 shows inter-annotator agreement between the two human teams, and between the human teams and GPT-4. Human and GPT-4 judgments match each other more than 98% of the time. Because of class imbalance (positive labels are rare, around 150 out of 8111), other measures of agreement such as Cohen’s $\kappa$ are in the mid 60’s, indicating substantial agreement between the human teams and between the human teams and GPT-4. Bootstrapped hypothesis tests find no significant difference between the human/human agreement rate and the human/GPT-4 agreement rate.
For the larger EvidenceBench-100k, over 150 million sentence-aspect pair judgements would need to be made. We switch the annotator to GPT4-o-mini-0718. We validated the quality of its annotation by re-running the hypothesis test on the same 50 randomly sampled papers, and found very similar Human & GPT agreement.
@@ 4 Experiment Setup
Evidence Retrieval @Optimal: denoted as ER@optimal. The smallest number of sentences to cover all aspects is denoted as Optimal. The average optimal number is only 4.6. The task is for a model to retrieve no more than the Optimal number of sentences to form the best evidence set relevant to the hypothesis. Evidence Retrieval @10 is denoted as ER@10.
Result Evidence Retrieval @Optimal: denoted as Result-ER@Optimal. The task restricts the model to retrieve no more than the optimal number of sentences, which is calculated by the minimum number of sentences required to cover all study aspects labeled as "Results". This labeling is done by GPT4-0125. Empirically, half of the aspects are labeled as "Results’, see Table 1. Result Evidence Retrieval @5 is denoted as Result-ER@5.
@@ 4.1 Evaluation strategies
Our evaluation metric is Aspect Recall, defined in Section 2.3.
@@ 5 Results
Section-by-Section Processing: On the other hand, from Table 4, Sec-by-Sec considerably improves Gemini and Claude’s performances, suggesting that the default longer-context version of the task hinders their retrieval abilities. Section-by-section is by far the most robust strategy observed here. We performed further analyses in Appendix J.
@@ 5.1 Fine-tuning and Evaluation on EvidenceBench-100k
Furthermore, we fine-tune two models: E5-v2 335M and Llama3-8B (sec-by-sec strategy) using the 80k training datapoints. We test them on the original EvidenceBench test set for the task of Result-ER@Optimal. We notice both fine-tuned models show significant improvements over their baselines as shown in Table 5(b). This shows the EvidenceBench-100k train set can be used for model developments. See full details of fine-tuning at Appendix H.
@@ 7 Conclusion
      Dataset License and Code LicenseA.1
@@
      Dataset Hosting, Accessibility and MaintenanceA.2
@@
      EvidenceBench Dataset StructureA.5
@@
      Guidelines for Hypothesis ValidationB.1
@@
      Annotation Guidelines for Alignment of Study Aspects and SentencesB.2
@@
        First Annotation StageB.2.1
@@
        Second Annotation StageB.2.2
@@
    Automated Alignment ProcedureC
@@
      Study Aspect DecompositionC.1
@@
    Experiment DetailsD
@@
      Default Prompt Template for Evidence RetrievalD.1
@@
      Prompt Template for Results Evidence Retrieval @Optimal and @5D.2
@@
      ICL Prompt for Evidence Retrieval @Optimal and @10D.3
@@
      Section-by-Section Prompt for Evidence Retrieval @Optimal and @10D.4
@@
      Instructions for Embedding ModelD.6
@@
      Standard Errors for Model EvaluationsD.7
  5. E.
    Model Sensitivity to Paraphrased HypothesisE
@@
    Hypothesis CollectionF
@@
    Data PreprocessingG
@@
      Harvesting Evidence SummaryG.1
@@
      Further Processing Evidence SummaryG.2
@@
      Identifying Suitable Evidence SummariesG.3
@@
      Human VerificationG.4
  8. H.
    Fine-tuning and Evaluation on EvidenceBench-100kH
@@
    Qualitative Analysis for GPT-4o on the Original EvidenceBenchI
@@
    Further AnalysesJ
@@ A.2 Dataset Hosting, Accessibility and Maintenance
A.3 A Motivating Example
@@
A.4 Dataset Collection and Processing
@@
A.5 EvidenceBench and EvidenceBench-100k Datasets Structure
@@ C.1 Study Aspect Decomposition
Appendix D Experiment Details
@@
D.1 Default Prompt Template for Evidence Retrieval
@@
D.2 Prompt Template for Results Evidence Retrieval @Optimal or @5
@@
D.3 ICL Prompt for Evidence Retrieval @Optimal and @10
@@
D.4 Section-by-Section Prompt for Evidence Retrieval @Optimal and @10
@@ D.5 Regeneration Prompt
D.6 Instructions for Embedding Model
@@
D.7 Standard Errors for Model Evaluations
@@
Appendix E Model Sensitivity to Paraphrased Hypothesis
@@
Appendix F Hypothesis Collection
@@
Appendix G Data Preprocessing
G.1 Harvesting Evidence Summary
@@
G.2 Further Processing Evidence Summary
@@
G.3 Identifying Suitable evidence summaries
@@ G.4 Human Verification
Appendix H Fine-tuning and Evaluation on EvidenceBench-100k
@@
Appendix I Qualitative Analysis for GPT-4o on the Original EvidenceBench
@@
Appendix J Further Analyses
@@
Section-level Reasoning is Sufficient: Table 4 shows that retrieving evidence section-by-section (Sec-by-Sec) achieves strong performance on the task. With Sec-by-Sec, a model can only read one section at a time, preventing it from reasoning across multiple sections. The strong performance of this method indicates that global reasoning across the entire research paper is not essential for retrieving evidence. This can be explained by the structure and organization of biomedical research papers, where the content of each section is relatively self-contained, and interaction across sections is sparse. This suggests that LLMs trained for much longer contexts (e.g., over 10,000 tokens) may not be necessary for this task.

### v2

@@ 1 Introduction
We conduct a benchmarking study on a variety of large language models (LLMs) and embedding models, which provide several insights. First, although LLMs still fall short of expert-level performance on this task, indicating that they cannot replace humans, they have the potential to assist them. Second, embedding models consistently underperform compared to large language models due to lack of global context from research papers. Third, we report that current LLMs trained for long document understanding still get "Lost in the Middle" (Liu et al., 2024). See Appendix K for details of our analyses.
@@ 2.1 Definitions
[…]
@@ 3.1 Data Sources
There are two data sources for EvidenceBench and EvidenceBench-100k. First, a collection of 107,887 CC-BY open-sourced biomedical research papers where each research paper represents a datapoint. Second, a collection of 44,772 review papers from PubMed Central. Each biomedical research paper has a corresponding evidence summary included in a review paper. Specifically, EvidenceBench has 426 datapoints and EvidenceBench-100k has 107,461 datapoints. Our datasets cover a wide range of biomedical topics, see Appendix A.3 for details.
@@ 3.2 Dataset Pipeline Overview
Harvesting the Evidence Summary: The highlighted portion in the review paper in Figure 2 is an evidence summary. Evidence summary has an XML citation embedded in it, so it can be identified and harvested by a deterministic algorithm. See Appendix H for details.
@@ 3.4 Alignment Annotation of Study Aspects and Sentences
Each team labeled 8111 (sentence, study aspect) pairs in total. Table 2 shows inter-annotator agreement between the two human teams, and between the human teams and GPT-4. Human and GPT-4 judgments match each other more than 98% of the time. Because of class imbalance (positive labels are rare, around 150 out of 8111), other measures of agreement such as Cohen’s $\kappa$ are in the mid 60’s, indicating substantial agreement between the human teams and between the human teams and GPT-4. Bootstrapped hypothesis tests find no significant difference between the human/human agreement rate and the human/GPT-4 agreement rate. To account for potential correlations among annotations from the same research paper, we additionally computed confidence intervals using hierarchical bootstrap sampling, and found similar results, shown in Appendix D.
For the larger EvidenceBench-100k, over 150 million sentence-aspect pair judgements would need to be made. We switch the annotator to GPT-4o-mini-0718. We validated the quality of its annotation by re-running the hypothesis test on the same 50 randomly sampled papers, and found very similar Human & GPT agreement. See Appendix D for 4o-mini’s results.
@@ 4 Experiment Setup
Evidence Retrieval @Optimal: denoted as ER@optimal. The smallest number of sentences to cover all aspects is denoted as Optimal. The average optimal number is only 4.6. The task is for a model to retrieve no more than the Optimal number of sentences to form the best evidence set relevant to the hypothesis. Evidence Retrieval @10 is denoted as ER@10. See Appendix E.1 for the prompt.
Result Evidence Retrieval @Optimal: denoted as Result-ER@Optimal. The task restricts the model to retrieve no more than the optimal number of sentences, which is calculated by the minimum number of sentences required to cover all study aspects labeled as "Results". This labeling is done by GPT4-0125. Empirically, half of the aspects are labeled as "Results’, see Table 1. Result Evidence Retrieval @5 is denoted as Result-ER@5. See Appendix E.2 for the prompt.
@@ 4.1 Evaluation strategies
Our evaluation metric is Aspect Recall, defined in Section 2.3. In Appendix E, we presented all prompts used in our experimental evaluation for LLMs, instructions for embedding models as well as standard errors for model evaluations.
@@ 5 Results
Section-by-Section Processing: On the other hand, from Table 4, Sec-by-Sec considerably improves Gemini and Claude’s performances, suggesting that the default longer-context version of the task hinders their retrieval abilities. Section-by-section is by far the most robust strategy observed here. We performed further analyses in Appendix K.
Reasoning-based LLMs: We tested two reasoning-based LLMs, Gemini 2.5 Pro (Google, 2025) and Claude 4 Sonnet (Anthropic, 2025), using the baseline prompt for ER@Optimal. As we see in Table 5, they scored 52.5 and 49.1 on ER@Optimal, respectively, numerically higher than GPT-4o’s performance. We leave a more extensive evaluation of reasoning models on EvidenceBench for future work.
@@
Table 5: Evaluation for reasoning-based LLMs using baseline prompt
Model Gemini 2.5 Pro Claude 4 Sonnet GPT-4o Claude 3 Opus Gemini 1.5
ER@Optimal 52.5 49.1 48.1 41.1 42.7
@@ 5.1 Fine-tuning and Evaluation on EvidenceBench-100k
Furthermore, we fine-tune two models: E5-v2 335M and Llama3-8B (sec-by-sec strategy) using the 80k training datapoints. We test them on the original EvidenceBench test set for the task of Result-ER@Optimal. We notice both fine-tuned models show significant improvements over their baselines as shown in Table 6(b). This shows the EvidenceBench-100k train set can be used for model developments. See full details of fine-tuning at Appendix I.
@@
5.2 Effect of Hypothesis Negation
In EvidenceBench, most of the sentences which are annotated as relevant to a hypothesis provide evidence in support of the hypothesis, rather than against it. To evaluate whether models are able to identify information which provides evidence against a hypothesis, we created a negated version of EvidenceBench by negating each hypothesis. See Appendix E.8 for the prompt and some examples of negated hypotheses. By negating each hypothesis, we have transformed the support relationships into refutation relationships.
Using the baseline prompt, we evaluated GPT-4o and Claude3-Opus on this negated version of the original EvidenceBench test set. In Table 4, the ER@Optimal for GPT-4o is 48.1 on the original set, the new experiment shows the ER@Optimal drops to 46.7 on the negated set. The ER@Optimal for Claude3 is 41.1 on the original set and the ER@Optimal drops to 36.7 on the negated set. The results demonstrate that both models show decreased performance when evaluating negated hypotheses, indicating that the task becomes more challenging when models must identify evidence that refutes rather than supports a claim. However, the decrease in performance is small for GPT-4o.
@@ 7 Conclusion
      Dataset License and Code License.A.1
@@
      Dataset Hosting, Accessibility and Maintenance.A.2
@@
      Dataset Collection and Processing.A.5
    6. 6.
      EvidenceBench Dataset Structure.A.6
@@
      Guidelines for Hypothesis Validation.B.1
@@
      Annotation Guidelines for Alignment of Study Aspects and Sentences.B.2
@@
        First Annotation Stage.B.2.1
@@
        Second Annotation Stage.B.2.2
@@
    Automated Alignment Procedure.C
@@
      Study Aspect Decomposition.C.1
@@
    Hypothesis Testing Results.D
  5. E.
    Experiment Details.E
@@
      Default Prompt Template for Evidence Retrieval.E.1
@@
      Prompt Template for Results Evidence Retrieval @Optimal and @5.E.2
@@
      ICL Prompt for Evidence Retrieval @Optimal and @10.E.3
@@
      Section-by-Section Prompt for Evidence Retrieval @Optimal and @10.E.4
@@
      Instructions for Embedding Model.E.6
@@
      Standard Errors for Model Evaluations.E.7
    8. 8.
      Negation Prompt.E.8
@@
    Model Sensitivity to Paraphrased Hypothesis.F
@@
    Hypothesis Collection.G
  8. H.
    Data Preprocessing.H
@@
      Harvesting Evidence Summary.H.1
@@
      Further Processing Evidence Summary.H.2
@@
      Identifying Suitable Evidence Summaries.H.3
@@
      Human Verification.H.4
@@
    Fine-tuning and Evaluation on EvidenceBench-100k.I
@@
    Qualitative Analysis for GPT-4o on the Original EvidenceBench.J
  11. K.
    Further Analyses.K
@@ A.2 Dataset Hosting, Accessibility and Maintenance
A.3 Dataset Coverage Analysis
In this paper, we introduce EvidenceBench with 426 papers and EvidenceBench-100k with 107,461 papers. Our two datasets are fully open-sourced under the CC-BY license and encompass a comprehensive range of biomedical topics. Specifically, we quantified topical diversity for the 100k papers using 2025 MeSH headings to demonstrate representativeness. The dataset covers all 16 top-level MeSH branches (A–N, V, Z), confirming coverage of every major biomedical domain. With 17,505 distinct MeSH descriptors representing 56.5% of the entire 2025 vocabulary (30,956 terms), the breadth matches that of a full-year MEDLINE slice. The branch-level distribution shows a Gini-Simpson diversity of 0.81 (where 1 indicates perfect evenness), demonstrating that no single area dominates the corpus. Additionally, MeSH tree numbers reach a median depth of 5 and 90th-percentile of 7, indicating the benchmark encompasses both fine-grained molecular and procedural topics alongside high-level concepts.
A.4 A Motivating Example
@@
A.5 Dataset Collection and Processing
@@
A.6 EvidenceBench and EvidenceBench-100k Datasets Structure
@@ C.1 Study Aspect Decomposition
Appendix D Hypothesis Testing Results
During hypothesis testing for automatic alignment annotation, although only 50 papers were annotated, each paper contains roughly 160 (sentence, aspect) pairs, totaling 8,111 pairs. In the paper, correlation and agreement metrics are calculated based on these 8,111 pairs and not on the 50 papers. Because annotations for sentences in the same paper are partially correlated, the effective sample size therefore lies between 50 and 8,111. In order to account for this correlation structure, we can use hierarchical bootstrap sampling (first sampling papers, then sampling sentences within papers) to compute confidence intervals and p-values. We provide results for GPT-4o and GPT-4o-mini on standard bootstrap sampling and hierarchical bootstrap sampling in Tables 7, 8, 9. 10.
Metrics Human & Human Average Human & GPT Average p-value
Exact Accuracy 98.8 $\pm$ 0.3 98.7 $\pm$ 0.2 0.21
F1 Binary 66.0 $\pm$ 6.5 64.6 $\pm$ 5.6 0.64
Cohen’s $\kappa$ 65.4 $\pm$ 6.6 63.9 $\pm$ 5.6 0.63
Spearman’s $\rho$ 65.4 $\pm$ 6.6 64.0 $\pm$ 5.6 0.65
Table 7: hypothesis testing result for GPT-4o on 8,111 pairs using standard bootstrap sampling
Metrics Human & Human Average Human & GPT Average p-value
Exact Accuracy 98.7 $\pm$ 0.5 98.6 $\pm$ 0.5 0.46
F1 Binary 66.5 $\pm$ 9.9 65.8 $\pm$ 8.8 0.91
Cohen’s $\kappa$ 65.8 $\pm$ 10.1 65.1 $\pm$ 9.0 0.90
Spearman’s $\rho$ 65.9 $\pm$ 10.0 65.3 $\pm$ 8.9 0.91
Table 8: Hypothesis testing result for GPT-4o using hierarchical bootstrap sampling.
Metrics Human & Human Average Human & GPT Average p-value
Exact Accuracy 98.7 $\pm$ 0.5 98.6 $\pm$ 0.5 0.134
F1 Binary 66.5 $\pm$ 12.3 64.0 $\pm$ 9.9 0.362
Cohen’s $\kappa$ 65.8 $\pm$ 12.6 63.3 $\pm$ 9.9 0.356
Spearman’s $\rho$ 65.9 $\pm$ 12.6 63.4 $\pm$ 9.9 0.360
Table 9: Hypothesis testing result for GPT-4o-mini on 8,111 pairs using standard bootstrap sampling.
Metrics Human & Human Average Human & GPT Average p-value
Exact Accuracy 98.7 $\pm$ 0.6 98.6 $\pm$ 0.6 0.339
F1 Binary 66.5 $\pm$ 9.9 64.0 $\pm$ 7.4 0.541
Cohen’s $\kappa$ 65.8 $\pm$ 10.1 63.3 $\pm$ 7.6 0.536
Spearman’s $\rho$ 65.9 $\pm$ 10.0 63.4 $\pm$ 7.6 0.533
Table 10: Hypothesis testing result for GPT-4o-mini using hierarchical bootstrap sampling.
Appendix E Experiment Details
@@
E.1 Default Prompt Template for Evidence Retrieval
@@
E.2 Prompt Template for Results Evidence Retrieval @Optimal or @5
@@
E.3 ICL Prompt for Evidence Retrieval @Optimal and @10
@@
E.4 Section-by-Section Prompt for Evidence Retrieval @Optimal and @10
@@ E.5 Regeneration Prompt
E.6 Instructions for Embedding Model
@@
E.7 Standard Errors for Model Evaluations
@@
E.8 Negation Prompt
We evaluate models on negated hypotheses using prompt in Figure 14. We give two example pairs of original hypothesis and negated hypothesis:
  • Original Hypothesis: Glucocorticoids cause cleft palate by inhibiting growth of the palatal shelves, preventing shelf contact and fusion.
    Negated Hypothesis: Glucocorticoids prevent cleft palate by promoting growth of the palatal shelves, facilitating shelf contact and fusion.
  • Original Hypothesis: The mortality rate from bladder cancer is higher among painters than in the general population.
    Negated Hypothesis: The mortality rate from bladder cancer is lower among painters than in the general population.
Figure 14: The prompt to negate a hypothesis
Appendix F Model Sensitivity to Paraphrased Hypothesis
@@
Appendix G Hypothesis Collection
@@
Appendix H Data Preprocessing
H.1 Harvesting Evidence Summary
@@
H.2 Further Processing Evidence Summary
@@
H.3 Identifying Suitable evidence summaries
@@ H.4 Human Verification
Appendix I Fine-tuning and Evaluation on EvidenceBench-100k
@@
Appendix J Qualitative Analysis for GPT-4o on the Original EvidenceBench
@@
Appendix K Further Analyses
@@
Section-level Reasoning is Sufficient: Table 4 shows that retrieving evidence section-by-section (Sec-by-Sec) achieves strong performance on the task. With Sec-by-Sec, a model can only read one section at a time, preventing it from reasoning across multiple sections. The strong performance of this method indicates that global reasoning across the entire research paper is not essential for retrieving evidence. This can be explained by the structure and organization of biomedical research papers, where the content of each section is relatively self-contained, and interaction across sections is sparse.
This suggests that LLMs trained for much longer contexts (e.g., over 10,000 tokens) may not be necessary for this task.